OpenStudy (anonymous):

0066 percent converted to decimal is

OpenStudy (anonymous):

is it 66 percent or 0.0066 percent?

OpenStudy (anonymous):

wait, i mean, i don't get ur question

OpenStudy (anonymous):

.0066

OpenStudy (anonymous):

00.66 is what I got

OpenStudy (anonymous):

0.0066% = 0.0066/100 = 0.00066/10 = 0.000066/1 = 0.000066 --> this is what i got:)